# Long-range angular correlations of particle displacements at a plastic-to-elastic transition in jammed amorphous solids

### v1

@@ Abstract
Understanding how a flow turns into an amorphous solid is a fundamental challenge in statistical physics, during which no apparent structural ordering appears. In the athermal limit, the two states are connected by a well-defined jamming transition, near which the solid is marginally stable. A recent mechanical response screening theory proposes an additional transition above jamming, called a plastic-to-elastic transition here, separating anomalous and quasi-elastic mechanical behavior. Through numerical inflation simulations in two dimensions, we show that the onsets of long-range radial and angular correlations of particle displacements decouple, occurring respectively at the jamming and plastic-to-elastic transitions. The latter is characterized by a power-law diverging correlation angle and a power-law spectrum of the displacements along a circle. This work establishes two-step transitions on the mechanical properties during “decompression melting” of an athermal over-jammed amorphous solid, reminiscent of the two-step structural melting of a crystal in two dimensions. In contradistinction with the latter, the plastic-to-elastic transition exists also in three dimensions.
@@
Introduction. When strained by an increasing deformation $\gamma$, a crystal displays a crossover from elastic to plastic responses. If the crystal is compressed, it responds elastically to a small $\gamma$ at any pressure $p$. The picture is dramatically changed in athermal amorphous solids near the jamming transition ($p\approx 0$). Such a solid is known to be marginally stable [1, 2, 3], with elasticity breaking down in the thermodynamic limit, even to infinitesimal mechanical perturbations [4, 5]. Indeed, the minimum strain $\gamma_{\rm min}$ required to trigger a plastic event vanishes in large systems [6, 7]. On the other hand, one expects elasticity restored in over-jammed solids well above the jamming transition ($p\gg 0$), when the inter-particle overlapping $\delta$ exceeds expected $\gamma_{\rm min}$, such that under small deformations the contact network remains unchanged. Thus there should be a plastic-to-elastic (PE) transition (or crossover) moving away from jamming.
@@
The expected PE transition has been recently explored by two approaches, both of which originated from theories. The first approach builds on the mean-field replica theory (MFRT) [3, 8], which predicts a Gardner transition in thermal hard spheres at $\varphi_{\rm G}^{<}$ below the jamming density $\varphi_{\rm J}$ [3, 9, 10, 11, 12], and another one in athermal soft spheres at $\varphi_{\rm G}^{>}$ above $\varphi_{\rm J}$ [5]. From the perspective of mechanical responses, the Gardner transition separates a stable phase where the system behaves like a standard elastic sold with a well-defined shear modulus $\mu$, and a marginally stable phase where $\mu$ is protocol-dependent [13]. From this viewpoint, the Gardner transition coincides with the PE transition, which occurs once the barriers between small meta-stable basins are no longer detectable when the system moves away from jamming. In unjammed thermal hard spheres, simulations have confirmed that the shear modulus $\mu_{\rm ZFC}$ measured by a zero-field cooled (ZFC) protocol and $\mu_{\rm FC}$ by a field-cooled (FC) protocol are identical ($\mu_{\rm ZFC}=\mu_{\rm FC}$) in the normal phase ($\varphi<\varphi_{\rm G}^{<}$), while $\mu_{\rm ZFC}>\mu_{\rm FC}$ in the Gardner phase ($\varphi_{\rm G}^{<}<\varphi<\varphi_{\rm J}$[14, 15]. In jammed athermal soft spheres, simulations report $\mu_{\rm ZFC}>\mu_{\rm FC}$ near jamming [16], but the transition $\varphi_{\rm G}^{>}$ itself has not been explored yet.
The second approach is motivated by a recent mechanical response screening theory (MRST) [17, 18, 19, 20, 21, 22]. The theory provides a possible PE transition in athermal over-jammed solids, between a quasi-elastic phase with quadrupole screening of the elastic field by plastic events, where plasticity simply renormalizes the elastic moduli, and an anomalous phase with dipole screening. This transition is observed in recent simulations, by showing the change of the screening parameter $\kappa$ from zero to a finite value at a non-zero pressure $p_{\rm c}$ [23]. However, several crucial questions remain to be answered regarding the nature of the transition: in particular, whether a sharp transition or a crossover occurs at $p_{\rm c}$ in the thermodynamic limit, and in the former case, whether it is associated with diverging long-range correlations.
@@
Figure 1: Radial displacement fields. Each column corresponds to typical response behavior at different pressures: (from right to left) $p=8.6\times 10^{2},5.2\times 10^{2},4.4\times 10^{-2}$ and 00, for $N=67352$. First row: heat maps of $\hat{d}_{r}(x,y)$ normalized by the maximum $\hat{d}_{r}^{\rm max}$. Second row: $\hat{d}_{r}(r)$, where the lines in (h,k) represent Eq. (1) with $d_{0}^{\rm eff}=0.054$ and $0.048$ respectively, and the line in (e) represents Eq. (2) with $d_{0}^{\rm eff}=0.072$ and $\kappa_{\rm fit}=0.026$. Third row: $\hat{d}_{r}(\theta)$ at $\hat{r}\equiv r/r_{\rm out}=0.15,0.5$, and $0.85$ (purple, green and red respectively).
@@
Following the second approach, here we show that the PE transition can be unambiguously differentiated from the jamming transition, although a finite-size scaling $p_{\rm c}\sim N^{-3/2}$ (for Hertzian interactions) suggests that $p_{\rm J}$ coincides with the jamming transition $p_{\rm c}=0$ in the thermodynamic limit. Physically, the jamming transition is associated with a diverging length due to long-range radial correlations in displacements, and, in contrast, the PE transition corresponds to a diverging angle due to long-range transverse correlations.
Three phases: quasi-elastic, plastic and unjammed.
@@
We simulate a bidisperse granular model of Hertzian (unless otherwise specified) particles in two dimensions, confined between a fixed outer boundary of radius $r_{\rm out}$ and an inner boundary of radius $r_{\rm in}$. A small particle, with a unit diameter ($R_{1}=0.5$) and a fixed position at the center ($r=0$), is inflated as $r_{\rm in}\to r_{\rm in}+d_{0}$, where $r_{\rm in}=R_{1}=0.5$ and $d_{0}=0.1$ (unless otherwise specified). The energy of the system is minimized after this instantaneous inflation, and the particle responses are characterized by the displacement field $\bm{\mathrm{d}}(r,\theta)$. In this study, we focus on the radial component of the displacement $d_{r}(r,\theta)$, where $d_{r}=\bm{\mathrm{d}}\cdot\hat{\bm{\mathrm{n}}}$ and $\hat{\bm{\mathrm{n}}}$ is the unit radial vector. The transverse component $d_{\theta}(r,\theta)$ is investigated in a separate study [24], revealing an interesting odd dipole screening effect. Details of the model and simulation methods are described in the Materials and Methods.
Figure 2: Two transitions: (a-b) jamming and (c-d) plastic-to-elastic. The characteristic length $\xi$ for the jamming transition as a function of (a) the packing density $\varphi$, and (b) $\varphi_{\rm J}-\varphi$, for a few different $N$. (c) The screen parameter $\kappa_{\rm fit}$ from fitting, and (d) the scaled pressure difference $\Delta p/p^{1/3}$ as functions of $p$, for $N=67352$. The blue data points (averaged over samples with $\kappa_{\rm fit}<0.01$ at the given $p$; error bars are invisible) have near-zero values, which can be unambiguously distinguished from others (red points; each point represents one sample). Note that the system is unjammed in (a-b) with $p=0$, and jammed in (c-d) with $p>0$.
@@
(i) A quasi-elastic solid phase at $p>p_{\rm c}$ and $\varphi>\varphi_{\rm c}$, where the mechanical response in the radial direction is elastic-like. The displacement field is dominated by dipolar and quadrupolar patterns (see Fig. 1g-l). Although the central inflation is isotropic, the actual displacement and force fields are generally anisotropic, revealing an important difference between a continuous medium and a discrete particle packing. In other words, the dipolar and quadrupolar patterns reflect disorder-induced breaking of rotational symmetry in granular models. Hexapoles and higher-order-poles exist but are originated from $1/f^{1.5}$ noise (see Fig. 3 and related discussions).
The displacement field $d_{r}(r,\theta)$ in this phase is consistent with the Michell solution that describes the mechanical response of a standard elastic medium [25]. Using the rescaled radial displacements $\hat{d}_{r}(r,\theta)=rd_{r}(r,\theta)$ (to remove the trivial $1/r$ decay of $d_{r}(r,\theta)$ at large $r$) and averaging out the angular dependence, the radial dependence of $\hat{d}_{r}(r)$ is expected to follow the solution of the normal elasticity theory,
@@
Eq. (1) is formally consistent with simulation results, with the inflation variable treated as a fitting parameter $d_{0}\to d_{0}^{\rm eff}$ (see Fig. 1h,k).
@@
Eq. (1) suggests that the response is elastic-like. The application of inflation may result in plastic events, which are typically quadrupolar in nature, aka Eshelby inclusions. At a large $p$, the plastic events are sparse, and the gradient of the quadrupole field $\bm{\mathrm{Q}}(r,\theta)$ is negligible. In this case, the effect of plasticity is to renormalize the elastic moduli, without modifying the overall behavior of the displacement field Eq. (1). For this reason, the high-pressure phase is termed a quasi-elastic phase. However, the angular dependence of $\hat{d}_{r}$ is not isotropic like in a true elastic medium. The oscillative behavior of $\hat{d}_{r}(\theta)$ (at a given $r$) suggests long-range correlations in the angular direction (Fig. 1i,l).
@@
(ii) A plastic solid phase at $0<p<p_{\rm c}$ and $\varphi_{\rm J}<\varphi<\varphi_{\rm c}$, where the system is overall jammed but the mechanical response anomalously disobeys elasticity. The displacement field $\hat{d}_{r}(r,\theta)$ can not be described by any regular patterns (see Fig. 1d-f). Under the isotropic assumption, the screening theory provides a solution of $\hat{d}_{r}(r)$ [17],
where $J_{1}$ and $Y_{1}$ are circular Bessel functions of the first and second kind respectively, and $\kappa$ is a screening parameter. Simulation data are fitted to Eq. (2) with two fitting parameters, $d_{0}^{\rm eff}$ and $\kappa_{\rm fit}$ (see Fig. 1e). Note that Eq. (1) is equivalent to setting $\kappa=0$ in Eq. (2).
The screening parameter $\kappa\sim 1/\ell_{\rm s}$ is the inverse of the screening length $\ell_{\rm s}$. The screening effect originates from an effective dipole field $\bm{\mathrm{P}}(r,\theta)$, which is the gradient of $\bm{\mathrm{Q}}(r,\theta)$ when the quadrupolar events are dense and non-uniformly distributed in the space [19]. While Eq. (2) gives rise to long-range correlations in the radial direction, $d_{r}(\theta)$ is noise-like, implying short-range correlations in the angular direction (Fig. 1f).
(iii) An unjammed phase at $p=0$ and $\varphi<\varphi_{\rm J}$, where the system is in a flow state. The influence of the central inflation propagates only up to a finite distance $\xi$, and $d_{r}(\theta)$ is noise-like (see Fig. 1a-c).
Let us highlight the essential differences between the three phases by summarizing the radial and angular dependence of $\hat{d}_{r}$ (see Fig. 1): (i) in the quasi-elastic solid phase, $\hat{d}_{r}(r)$ follows the elasticity theory Eq. (1) and $\hat{d}_{r}(\theta)$ is oscillative (both long-ranged); (ii) in the plastic solid phase, $\hat{d}_{r}(r)$ follows the MRST Eq. (2) and $\hat{d}_{r}(\theta)$ is noise-like (long-ranged in $r$ and short-ranged in $\theta$); (iii) in the unjammed phase, $\hat{d}_{r}(r)$ decays to zero at a finite $\xi$ and $\hat{d}_{r}(\theta)$ is noise-like (both short-ranged). Below we characterize these differences in detail, focusing on the features of the two transitions between the three phases.
Two transitions: jamming and plastic-to-elastic. The jamming transition between unjammed and plastic phases has been extensively investigated in previous studies [26, 27, 28, 29, 30, 31, 1, 32]. In Ref. [29], the radial correlations between particles’ non-affine displacements are measured in athermal quasi-static shear simulations with Lees-Edwards boundary conditions, giving a diverging correlation length $\xi$ when the jamming transition is approached from below,
$\xi\sim(\varphi_{\rm J}-\varphi)^{-\nu},$ (3)
with $\nu\sim 0.8-1.0$. In this study, we estimate $\xi$ by the length scale at which the radial displacement $d_{r}(r)$ decays to a small threshold value, $d_{r}(r=\xi)=10^{-3}$ (Fig. 2a). The power-law divergence of $\xi$ is confirmed in Fig. 2b, with $\varphi_{\rm J}=0.842$ [33] and $\nu\approx 0.8$ [29] consistent with previous studies.
@@
The PE transition between plastic and quasi-elastic phases was only noticed very recently [23]. To provide a rough estimation of the PE transition pressure $p_{\rm c}$, we follow the strategy in  [23]: the data of $\hat{d}_{r}(r)$ at different $p$ are fitted to Eq. (2), and the fitting parameter $\kappa_{\rm fit}$ is plotted as a function of $p$ in Fig. 2c. At high $p$, $\kappa_{\rm fit}\approx 0$ (blue points), consistent with Eq. (1). The non-zero $\kappa_{\rm fit}$ appears only below $p_{\rm c}\approx 20$ for the system of $N=67352$ particles (red points). In Ref. [23], a selection principle of $\kappa$ is proposed: the $\hat{d}_{r}(r)$ curves measured in simulations preferably select a series of $\kappa_{n}\approx n\pi/r_{\rm out}$ with a decreasing probability when $n$ is increased, where $n=1,2,\ldots$ At $\kappa=\kappa_{n}$ the screening effects are significantly strong, which can be seen from the behavior of Eq. (2), since the amplitude $\hat{d}_{\rm amp}(\kappa=\kappa_{n})$ diverges. Here $\hat{d}_{\rm amp}(\kappa)$ is the extreme value of the $\hat{d}_{r}(r)$ curve for the given $\kappa$. In Fig. 2c, the first two $\kappa_{n}$, i.e., $\kappa_{1}\approx 0.022$ and $\kappa_{1}\approx 0.041$, are indicated by dashed lines. The data points of non-zero $\kappa_{\rm fit}$ all fall in the range between $\kappa_{1}$ and $\kappa_{2}$. The clear gap between $\kappa=0$ and $\kappa_{1}$ supports the selection principle. Meanwhile, we observe a wide fluctuation of $\kappa_{\rm fit}$ at $p<p_{\rm c}$, possibly due to strong sample-to-sample fluctuations in finite-sized systems and the uncertainty in the fitting.
To examine the reliability of $p_{\rm c}$ estimated above, in this study we propose two additional methods to determine $p_{\rm c}$. In Fig. 2d, we plot the pressure difference $\Delta p\equiv p_{\rm final}-p$ between the pressures after and before the central inflation (both pressures are measured with the energy minimized). To properly normalize the pressure, note that $p\sim F\sim\delta^{3/2}$, where the second relation between the inter-particle force $F$ and the inter-particle linear overlap $\delta$ holds for the Hertzian potential. From this, one obtains $d\delta\sim dp/p^{1/3}$. The data (red points) in Fig. 2d shows that $\Delta p/p^{1/3}$ fluctuates around $-0.05$ below $p_{\rm c}\approx 20$. The fluctuation is large, but nevertheless the negative values confirm the anomaly, compared to the small positive values expected from the elastic behavior (blue points). Note that this second approach can be straightforwardly accessed in granular experiments [18]. In the next section, we will discuss a third way to determine $p_{\rm c}$, which gives the most accurate estimate.
The above analysis shows that jamming and PE transitions are two separate transitions, at least in finite-sized systems. For a finite $N$, the PE transition occurs at a finite $p_{\rm c}$, while the jamming transition occurs at $p=0$ for any $N$. A finite-size analysis will be presented below to extrapolate the asymptotic behavior in the thermodynamic limit.
Emergence of long-range angular correlations at the plastic-to-elastic transition. As discussed above, the onset of long-range displacement correlations in the radial direction occurs at the jamming transition $p=p_{\rm J}=0$ and $\varphi=\varphi_{\rm J}$ (Fig. 2a-b). Next we show that the onset of long-range displacement correlations in the angular direction appears differently at the PE transition $p=p_{\rm c}$. To do that, we define an angular correlation function of the radial displacements,
@@
Fig. 3a reports $G(\theta)$ for $\hat{r}=0.5$ at different pressures. In the large pressure limit, the displacement field is dominated by dipolar patterns as shown in Fig. 1j, which corresponds to $G(\theta)=\cos(\theta)$. With decreasing $p$, quadrupolar and other patterns (see Fig. 1g) appear, and $G(\theta)$ deviates from the cosine function. Below $p_{\rm c}$, $G(\theta)$ rapidly decays to zero, indicating short-range correlations. To quantify the behavior change of $G(\theta)$, we define an order parameter $G^{*}\equiv-G(\theta=\pi)$, and a correlation angle $G(\theta=\theta^{*})=G_{\rm th}$ with a threshold $G_{\rm th}=0.1$. Fig. 4a-b show that both $G^{*}$ and $\theta^{*}$ grow with $p$.

### v2

@@ Abstract
Understanding how a fluid turns into an amorphous solid is a fundamental challenge in statistical physics, during which no apparent structural ordering appears. In the athermal limit, the two states are connected by a well-defined jamming transition, near which the solid is marginally stable. A recent mechanical response screening theory proposes an additional transition above jamming, called a plastic-to-elastic transition here, separating anomalous and quasielastic mechanical behavior. Through numerical inflation simulations in two dimensions, we show that the onsets of long-range radial and angular correlations of particle displacements decouple, occurring, respectively, at the jamming and plastic-to-elastic transitions. The latter is characterized by a power-law diverging correlation angle and a power-law spectrum of the displacements along a circle. This work establishes two-step transitions on the mechanical properties during “decompression melting” of an athermal overjammed amorphous solid, reminiscent of the two-step structural melting of a crystal in two dimensions. In contradistinction with the latter, the plastic-to-elastic transition exists also in three dimensions.
preprint: APS/123-QED
@@
Introduction. When strained by an increasing deformation $\gamma$, a crystal displays a crossover from elastic to plastic responses. If the crystal is compressed, it responds elastically to a small $\gamma$ at any pressure $p$. The picture is dramatically changed in athermal amorphous solids near the jamming transition ($p\approx 0$). Such a solid is known to be marginally stable [1, 2, 3], with elasticity breaking down in the thermodynamic limit, even to infinitesimal mechanical perturbations [4, 5]. Indeed, the minimum strain $\gamma_{\rm min}$ required to trigger a plastic event vanishes in large systems [6, 7]. On the other hand, one expects elasticity restored in overjammed solids well above the jamming transition ($p\gg 0$), when the interparticle overlapping $\delta$ exceeds expected $\gamma_{\rm min}$, such that under small deformations the system can be effectively considered as an elastic medium. Thus there should be a plastic-to-elastic (PE) transition (or crossover) moving away from jamming.
@@
The expected PE transition has been recently explored by two approaches, both of which originated from theories. The first approach builds on the mean-field replica theory (MFRT) [3, 8], which predicts a Gardner transition [3, 9, 10, 11, 12] in athermal soft spheres at $\varphi_{\rm G}^{>}$ above the jamming density $\varphi_{\rm J}$ [5]. The Gardner transition separates the elastic phase and the plastic phase where the shear modulus is protocol dependent due to marginal stability [13, 14, 15, 16].
@@
Figure 1: Radial displacement fields. Each row corresponds to typical response behavior at different pressures (from top to bottom): $p=0,4.4\times 10^{-2},5.2\times 10^{2},8.6\times 10^{2}$, for $N=67352$. First column: heat maps of $\hat{d}_{r}(x,y)$ normalized by the maximum $\hat{d}_{r}^{\rm max}$. Second column: $\hat{d}_{r}(r)$, where the lines in (h,k) represent Eq. (1) with $d_{0}^{\rm eff}=0.054$ and 0.048, respectively, and the line in (e) represents Eq. (2) with $d_{0}^{\rm eff}=0.072$ and $\kappa_{\rm fit}=0.026$. Third column: $\hat{d}_{r}(\theta)$ at $\hat{r}\equiv r/r_{\rm out}=0.15,0.5$, and $0.85$ (purple, green, and red, respectively).
The second approach is motivated by a recent mechanical response screening theory (MRST) [17, 18, 19, 20, 21, 22]. The theory provides a possible PE transition in athermal overjammed solids, between a quasielastic phase with quadrupole screening of the elastic field by plastic events, where plasticity simply renormalizes the elastic moduli, and an anomalous phase with dipole screening. This transition is observed in recent simulations, by showing the change of the screening parameter $\kappa$ from zero to a finite value at a nonzero pressure $p_{\rm c}$ [23]. However, several crucial questions remain to be answered regarding the nature of the transition: in particular, whether a sharp transition or a crossover occurs at $p_{\rm c}$ in the thermodynamic limit, and in the former case, whether it is associated with diverging long-range correlations.
@@
Following the second approach, here we show that the PE transition can be unambiguously differentiated from the jamming transition: the jamming transition is associated with a diverging length due to long-range radial correlations in displacements, and, in contrast, the PE transition corresponds to a diverging angle due to long-range transverse correlations. It should be noted that the PE transition captures the change of the macroscopic mechanical response behavior, which is beyond the first contact change events at the microscopic level [24, 25].
@@
Three phases: quasielastic, plastic and unjammed. We simulate a bidisperse granular model of Hertzian (unless otherwise specified) particles in two dimensions, confined between a fixed outer boundary of radius $r_{\rm out}$ and an inner boundary of radius $r_{\rm in}$. A small particle, with a unit diameter ($R_{1}=0.5$) and a fixed position at the center ($r=0$), is inflated as $r_{\rm in}\to r_{\rm in}+d_{0}$, where $r_{\rm in}=R_{1}=0.5$ and $d_{0}=0.1$ (unless otherwise specified). The energy of the system is minimized after this instantaneous inflation, and the particle responses are characterized by the displacement field $\bm{\mathrm{d}}(r,\theta)$. In this study, we focus on the radial component of the displacement $d_{r}(r,\theta)$, where $d_{r}=\bm{\mathrm{d}}\cdot\hat{\bm{\mathrm{n}}}$ and $\hat{\bm{\mathrm{n}}}$ is the unit radial vector. The transverse component $d_{\theta}(r,\theta)$ is investigated in a separate study [26], revealing an interesting odd dipole screening effect. Details of the model and simulation methods are described in Appendixes A and B.
@@
(i) A quasielastic solid phase at $p>p_{\rm c}$ and $\varphi>\varphi_{\rm c}$, where the mechanical response in the radial direction is elasticlike. The displacement field is dominated by dipolar and quadrupolar patterns [see Figs. 1(g)-1(l)]. Although the central inflation is isotropic, the actual displacement and force fields are generally anisotropic. Hexapoles and higher-order poles exist but are originated from $1/f^{1.5}$ noise (see Fig. 3 and related discussions).
The displacement field $d_{r}(r,\theta)$ in this phase is consistent with the Michell solution that describes the mechanical response of a standard elastic medium [27]. Using the rescaled radial displacements $\hat{d}_{r}(r,\theta)=rd_{r}(r,\theta)$ to remove the trivial $1/r$ decay of $d_{r}(r,\theta)$ at large $r$ and averaging out the angular dependence, the radial dependence of $\hat{d}_{r}(r)$ is expected to follow the solution of the normal elasticity theory,
@@
Eq. (1) is consistent with simulation results, with the inflation variable treated as a fitting parameter $d_{0}\to d_{0}^{\rm eff}$ [see Figs. 1(h) and 1(k)].
@@
Eq. (1) suggests that the response is elasticlike. The application of inflation may result in plastic events, which are typically quadrupolar in nature, aka Eshelby inclusions. At a large $p$, the plastic events are sparse, and the gradient of the quadrupole field $\bm{\mathrm{Q}}(r,\theta)$ is negligible. In this case, the effect of plasticity is to renormalize the elastic moduli, without modifying the overall behavior of the displacement field Eq. (1). For this reason, the high-pressure phase is termed a quasielastic phase. However, the angular dependence of $\hat{d}_{r}$ is not isotropiclike in a true elastic medium (see Sec. S1 in Supplemental Material (SM) [28]). The oscillative behavior of $\hat{d}_{r}(\theta)$ (at a given $r$) suggests long-range correlations in the angular direction [Figs. 1(i) and 1(l)).
(ii) A plastic solid phase at $0<p<p_{\rm c}$ and $\varphi_{\rm J}<\varphi<\varphi_{\rm c}$, where the system is overall jammed but the mechanical response anomalously disobeys elasticity. The displacement field $\hat{d}_{r}(r,\theta)$ can not be described by any regular patterns [see Figs. 1(d)-(f)]. Under the isotropic assumption, the screening theory provides a solution of $\hat{d}_{r}(r)$ [17],
where $J_{1}$ and $Y_{1}$ are circular Bessel functions of the first and second kind, respectively, and $\kappa$ is a screening parameter. Simulation data are fitted to Eq. (2) with two fitting parameters, $d_{0}^{\rm eff}$ and $\kappa_{\rm fit}$ [see Fig. 1(e)]. Note that Eq. (1) is equivalent to setting $\kappa=0$ in Eq. (2). See Sec. S2 in SM [28, 29, 30] for ensemble-averaged data.
The screening parameter $\kappa\sim 1/\ell_{\rm s}$ is the inverse of the screening length $\ell_{\rm s}$. The screening effect originates from an effective dipole field $\bm{\mathrm{P}}(r,\theta)$, which is the gradient of $\bm{\mathrm{Q}}(r,\theta)$ when the quadrupolar events are dense and non-uniformly distributed in the space [19]. While Eq. (2) gives rise to long-range correlations in the radial direction, $d_{r}(\theta)$ is noiselike, implying short-range correlations in the angular direction [Fig. 1(f)].
(iii) An unjammed phase at $p=0$ and $\varphi<\varphi_{\rm J}$, where the system is in a fluid state. The influence of the central inflation propagates only up to a finite distance $\xi$, and $d_{r}(\theta)$ is noiselike [see Figs. 1(a)-1(c)].
Figure 2: Plastic-to-elastic transition. (a) The screening parameter $\kappa_{\rm fit}$ from fitting, and (b) the scaled pressure difference $\Delta p/p^{1/3}$ as functions of $p$, for $N=67352$. The blue data points (averaged over samples with $\kappa_{\rm fit}<0.01$ at the given $p$; error bars are invisible) have near-zero values, which can be unambiguously distinguished from others (red points; each point represents one sample).
Let us highlight the essential differences between the three phases by summarizing the radial and angular dependence of $\hat{d}_{r}$ (see Fig. 1): (i) in the quasielastic solid phase, $\hat{d}_{r}(r)$ follows the elasticity theory Eq. (1) and $\hat{d}_{r}(\theta)$ is oscillative (both long ranged); (ii) in the plastic solid phase, $\hat{d}_{r}(r)$ follows the MRST Eq. (2) and $\hat{d}_{r}(\theta)$ is noiselike (long ranged in $r$ and short ranged in $\theta$); (iii) in the unjammed phase, $\hat{d}_{r}(r)$ decays to zero at a finite $\xi$ and $\hat{d}_{r}(\theta)$ is noiselike (both short ranged).
Two transitions: jamming and plastic-to-elastic. The jamming transition between unjammed and plastic phases has been extensively investigated in previous studies [31, 32, 33, 34, 35, 36, 1, 37]. In Ref. [34], the radial correlations between particles’ nonaffine displacements are measured in athermal quasistatic shear simulations with Lees-Edwards boundary conditions, giving a diverging correlation length $\xi$ when the jamming transition is approached from below,
$\xi\sim(\varphi_{\rm J}-\varphi)^{-\nu},$ (3)
with $\nu\sim 0.8-1.0$. In this study, we estimate $\xi$ by the length scale at which the radial displacement $d_{r}(r)$ decays to a small threshold value, $d_{r}(r=\xi)=10^{-3}$ (see Sec. S3 in SM [28]). The power-law divergence of $\xi$ is confirmed with $\varphi_{\rm J}=0.842$ [38] and $\nu\approx 0.8$ [34] consistent with previous studies.
The PE transition between plastic and quasielastic phases was only noticed very recently [23]. To provide a rough estimation of the PE transition pressure $p_{\rm c}$, we follow the strategy in  [23]: the data of $\hat{d}_{r}(r)$ at different $p$ are fitted to Eq. (2), and the fitting parameter $\kappa_{\rm fit}$ is plotted as a function of $p$ in Fig. 2(a). At high $p$, $\kappa_{\rm fit}\approx 0$ (blue points), consistent with Eq. (1). The nonzero $\kappa_{\rm fit}$ appears only below $p_{\rm c}\approx 20$ for the system of $N=67352$ particles (red points). In Ref. [23], a selection principle of $\kappa$ is proposed: the $\hat{d}_{r}(r)$ curves measured in simulations preferably select a series of $\kappa_{n}\approx n\pi/r_{\rm out}$ with a decreasing probability when $n$ is increased, where $n=1,2,\ldots$ At $\kappa=\kappa_{n}$ the screening effects are significantly strong, which can be seen from the behavior of Eq. (2), since the amplitude $\hat{d}_{\rm amp}(\kappa=\kappa_{n})$ diverges. Here $\hat{d}_{\rm amp}(\kappa)$ is the extreme value of the $\hat{d}_{r}(r)$ curve for the given $\kappa$. In Fig. 2(a), the first two $\kappa_{n}$, i.e., $\kappa_{1}\approx 0.022$ and $\kappa_{2}\approx 0.041$, are indicated by dashed lines. The data points of nonzero $\kappa_{\rm fit}$ all fall in the range between $\kappa_{1}$ and $\kappa_{2}$. The clear gap between $\kappa=0$ and $\kappa_{1}$ supports the selection principle. Meanwhile, we observe a wide fluctuation of $\kappa_{\rm fit}$ at $p<p_{\rm c}$, possibly due to strong sample-to-sample fluctuations in finite-sized systems and the uncertainty in the fitting.
@@
To examine the reliability of $p_{\rm c}$ estimated above, in this study we propose two additional methods to determine $p_{\rm c}$. In Fig. 2(b), we plot the pressure difference $\Delta p\equiv p_{\rm final}-p$ between the pressures after and before the central inflation (both pressures are measured with the energy minimized). To properly normalize the pressure, note that $p\sim F\sim\delta^{3/2}$, where the second relation between the interparticle force $F$ and the interparticle linear overlap $\delta$ holds for the Hertzian potential. From this, one obtains $d\delta\sim dp/p^{1/3}$. The data (red points) in Fig. 2(b) shows that $\Delta p/p^{1/3}$ fluctuates around $-0.05$ below $p_{\rm c}\approx 20$. The fluctuation is large, but nevertheless the negative values confirm the anomaly, compared to the small positive values expected from the elastic behavior (blue points). Note that this second approach can be straightforwardly accessed in granular experiments [18]. In the next section, we will discuss a third way to determine $p_{\rm c}$, which gives the most accurate estimate.
@@
Emergence of long-range angular correlations at the plastic-to-elastic transition. As discussed above, the onset of long-range displacement correlations in the radial direction occurs at the jamming transition $p=p_{\rm J}=0$ and $\varphi=\varphi_{\rm J}$. Next we show that the onset of long-range displacement correlations in the angular direction appears at the PE transition $p=p_{\rm c}$. We define an angular correlation function of the radial displacements,
@@
Figure  3(a) reports $G(\theta)$ for $\hat{r}=0.5$ at different pressures. In the large pressure limit, the displacement field is dominated by dipolar patterns as shown in Fig. 1(j), which corresponds to $G(\theta)=\cos(\theta)$. With decreasing $p$, quadrupolar and other patterns [see Fig. 1(g)] appear, and $G(\theta)$ deviates from the cosine function. Below $p_{\rm c}$, $G(\theta)$ rapidly decays to zero, indicating short-range correlations. To quantify the behavior change of $G(\theta)$, we define an order parameter $G^{*}\equiv-G(\theta=\pi)$, and a correlation angle $G(\theta=\theta^{*})=G_{\rm th}$ with a threshold $G_{\rm th}=0.1$ (see Sec. S4 in SM [28]). Figures  4(a) and 4(b) show that both $G^{*}$ and $\theta^{*}$ grow with $p$.